Solve the given differential equation by using an appropriate substitution. The DE is a Bernoulli equation.



x dy/dx + y = 1/y^2

1
Expert's answer
2021-11-09T16:28:50-0500

divide left and right sides of equation by "-(1-y^3)\/y^2" :


"\frac{y^2}{y^3-1}dy=-\frac{dx}{x}"


"\int\frac{y^2}{y^3-1}dy=-\int\frac{dx}{x}"


"\frac{ln(y^3-1)}{3}=c-lnx"


"y=\sqrt[3]{\frac{c_1}{x^3}+1}"